OpenStudy (anonymous):

Vivian combined different amounts of white paint, blue paint, and green paint to make 145 milliliters (ml) of paint of a new shade. The amount of white paint was x ml. The amount of blue paint was 10 ml less than twice the amount of white paint, and the amount of green paint was half the amount of blue paint. This is represented in the equation below. x + (2x – 10) + 1/2 (2x – 10) = 145 What was the difference in the amounts of white paint and blue paint Vivian combined? 30 ml 35 ml 40 ml 45 ml

OpenStudy (anonymous):

" x " - amount of white paint "2x-10" - amount of green paint "1/2 ( 2x -10)" - amount of blue paint

OpenStudy (anonymous):

difference means subtracting one from the other

jimthompson5910 (jim_thompson5910):

x + (2x – 10) + 1/2 (2x – 10) = 145 x + (2x – 10) + 1/2 (2x) – 1/2(10) = 145 x + 2x – 10 + x – 5 = 145 4x – 15 = 145 I'll let you finish solving for x

OpenStudy (anonymous):

@jim_thompson5910 is solving for x really necessary?

jimthompson5910 (jim_thompson5910):

yes, you need to know how much white paint there is

OpenStudy (anonymous):

if blue paint is x-5 and white paint is x then the difference would be ...

jimthompson5910 (jim_thompson5910):

btw, it says "The amount of blue paint was 10 ml less than twice the amount of white paint", so amount of blue paint: 2x - 10

jimthompson5910 (jim_thompson5910):

you mixed up blue and green

OpenStudy (anonymous):

ah i see crap

jimthompson5910 (jim_thompson5910):

no biggie, you just got a free color out of it: yellow

OpenStudy (anonymous):

@jim_thompson5910 I did something wrong i got 15 :/

OpenStudy (anonymous):

wait no its 40 right?@jim_thompson5910

jimthompson5910 (jim_thompson5910):

4x – 15 = 145 4x = 145+15 4x = 160 x = 160/4 x = 40 so you have 40 mL of white paint amount of blue paint: 2x - 10 amount of blue paint: 2*40 - 10 amount of blue paint: 60 - 10 amount of blue paint: 50 ------------------------- Difference: (amount of blue) - (amount of white) Difference: (50) - (40) Difference: 10 hmm but 10 isn't an answer choice

jimthompson5910 (jim_thompson5910):

oops typo lol one sec

jimthompson5910 (jim_thompson5910):

amount of blue paint: 2x - 10 amount of blue paint: 2*40 - 10 amount of blue paint: 80 - 10 amount of blue paint: 70 --------------------------------- Difference: (amount of blue) - (amount of white) Difference: (70) - (40) Difference: 30

jimthompson5910 (jim_thompson5910):

there, fixed

OpenStudy (anonymous):

oh okay i get it thank you (: